I need help on this equation:
y=[x+3]-4=

I need the x-intercepts(the negative and positive).....im stuck at that part..
Thanks

Answer by rapaljer(4551)

After I solved the following exercise, it occurred to me that you MAY have meant this to be an absolute value problem!! Did you intentionally use brackets [ ] in the problem? Brackets in a math problem do NOT mean absolute value. Before continuing, let's make sure you MEANT to give me a bracket function. If you meant absolute value, then IGNORE all of this HARD work that I did trying to solve a BRACKET FUNCTION!!

Bracket function, the greatest integer function!! What this means is that whatever you get when you substitute a value of x, you round it down to the greatest integer.

y= [x+3] -4 means that
if x= 0, then y= [0+3] - 4 = 3 - 4 = -1
if x=.5, then y= [.5+3] - 4 = 3 - 4 = -1
if x=1, then y = [1+3] - 4 = 4 - 4 = 0 which is an x-intercept
if x=1.5, then y = [1.5+3] - 4 = 4 - 4 = 0 which is also an x-intercept
if x=1.9, then y = [1.9+3] - 4 = 4 - 4 = 0 which is also an x-intercept
if x=2, then y = [2+3] - 4 = 5 - 4 = 1

In other words, everything between 1 and 2 including the 1 but NOT including the 2 will be x-intercepts. This is called the "unit step function. However, this may NOT have been what you were asking!

R^2 at SCC